if coefficiet of a quadratic equation is in A.P. Then find sum of alpha beta + product of alpha beta

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
Assume ax2+bx+c be the quadratic equation a,b,c being in AP so 2b=a+c
and p,q be the roots of equation 
We need p+q+pq
So p+q =-b/a
pq=c/a So 
p+q+pq= (c-b)/a
c-b=b-a
so 
p+q+pq=b/a-1